# SP-Chain: Boosting Intra-Shard and Cross-Shard Security and Performance in Blockchain Sharding

### v1

@@ Abstract
A promising way to overcome the scalability limitations of the current blockchain is to use sharding, which is to split the transaction processing among multiple, smaller groups of nodes. A well-performed blockchain sharding system requires both high performance and high security in both intra- and cross-shard perspectives. However, existing protocols either have issues on protecting security or trade off great performance for security. In this paper, we propose SP-Chain, a blockchain sharding system with enhanced Security and Performance for both intra- and cross-shard perspectives. For intra-shard aspect, we design a two-phase concurrent voting scheme to provide high system throughput and low transaction confirmation latency. Moreover, we propose an efficient unbiased leader rotation scheme to ensure high performance under malicious behavior. For cross-shard aspect, a proof-assisted efficient cross-shard transaction processing mechanism is proposed to guard the cross-shard transactions with low overhead. We implement SP-Chain based on Harmony, and evaluate its performance via large-scale deployment. Extensive evaluations suggest that SP-Chain can process more than 10,000 tx/sec under malicious behaviors with a confirmation latency of 7.6s in a network of 4,000 nodes.
@@
Blockchain, blockchain sharding, intra-shard consensus, cross-shard transaction processing
1 Introduction
Since the advent of Bitcoin [27], blockchain systems have continued to have a significant impact on society. However, the low system throughput and high transaction confirmation delays of such systems greatly hinder their usability across various infrastructures and applications. Consequently, sharding [8, 23], a promising blockchain scaling solution, has been proposed. Herein, the entire blockchain state is divided into multiple non-overlapping shards, each maintained by a group of nodes.
@@ 1 Introduction
Intra-shard security and performance: Within each shard, a specific leader is typically selected to propose blocks. Malicious leaders or attacks on leaders can severely affect intra-shard security and performance. Specifically, most existing sharding protocols [23, 17, 38, 12, 18] require each shard’s leader to produce blocks within a certain time frame and to know the next leader in advance. This allows attackers to easily launch targeted attacks on the leaders [20], weakening system security (for example, colluding with the leader for Byzantine behavior or launching DDoS attacks against the leader). Moreover, when a leader is found to be malicious or under attack, a complex view change process is required to elect a new leader [5], thereby reducing system performance.
Another issue is that many existing sharding protocols adopt generic Byzantine Fault Tolerance (BFT) consensus protocols (such as PBFT) [6] for intra-shard consensus, requiring extensive communication among shard members. Specifically, BFT-type consensus protocols typically require multi-stage communication (such as pre-preparation, preparation, commit) to ensure security in poor network connectivity and high latency situations. However, these protocols are inefficient when applied to blockchain sharding. In sharding systems, each shard generally has high network connectivity and a low and fixed message propagation cap [37, 17, 13, 38, 21]. Under such conditions, a more efficient consensus protocol with less communication overhead can be proposed for sharding systems.
@@
Random Leader Rotation. To address the first challenge, the leader of each shard is frequently, randomly, and automatically changed. Specifically, leaders are rotated with each block as the cycle. To prevent attackers from knowing subsequent leaders in advance, leaders are selected based on distributed randomness only before each block is proposed. To ensure performance and security during the distributed randomness generation process, we propose a chain-based randomness generation mechanism. Herein, nodes within each shard use the signature information of previously confirmed blocks to efficiently generate unbiased distributed randomness. In addition, the automatic leader rotation mechanism eliminates the originally complex view change process, further improving the efficiency of leader election.
@@
To further improve consensus efficiency, we parallelize steps that are executed serially in traditional consensus protocols. In traditional protocols, the leader first packages transactions into a block. Then, the leader broadcasts the block. Nodes receiving the block will vote and reach a consensus. After reaching a consensus, nodes insert the block into the blockchain. Our concurrent voting protocol parallelizes these steps. Specifically, when a shard’s leader produces a new block, the members of that shard vote on and reach a consensus on the previous block. Also, while the leader broadcasts the new block, each node inserts the previous block into the blockchain. This proposed scheme provides fast block generation and confirmation rates, bringing high throughput and low latency to the system.
@@
In response to the above challenges, we propose SP-Chain, a secure and efficient blockchain sharding system that enhances both intra-shard and inter-shard security and performance. Specifically, the leader rotation protocol provides a secure and efficient intra-shard leader election, the two-phase concurrent voting consensus protocol improves intra-shard consensus efficiency, and the proof-assisted efficient cross-shard transaction processing ensures cross-shard transaction security without excessive communication overhead.
We implemented the prototype of SP-Chain based on Harmony [1], a well-known public blockchain sharding project that was ranked among the top 50 in the cryptocurrency space in terms of market capitalization. We conducted extensive experiments with different numbers of shards on Amazon EC2. The experimental results show that SP-Chain can still maintain high efficiency under malicious behavior. In a network with 4,000 nodes, its throughput exceeds 10,000 tx/sec, and the transaction confirmation delay is 7.6 seconds.
2 Background and Related Work
2.1 Blockchain Sharding
@@
A well-performed blockchain sharding system needs to ensure good security as well as high performance. Unlike traditional blockchain systems, in blockchain sharding systems, there are a large number of cross-shard transactions that are transmitted among shards [33, 17, 37, 28, 14, 19, 29]. Nodes need to not only process transactions within a shard, but also handle transactions across shards. Therefore, it is neccesary to consider both intra- and cross-shard aspects when analyzing the security and performance of blockchain sharding. However, previous works have problems with both intra- and cross-shard security and performance.
Issues from Intra-Shard Aspect. Existing sharding works reach consensus with limited efficiency. A major reason is that most of them [23, 17, 13, 9, 38, 12, 14, 18, 22] use BFT-typed consensus protocols suitable for scenarios with poor network conditions. Those protocols usually require multiple rounds of communication ($\geq$3) to reach consensus, slowing down the efficiency of reaching consensus. However, it is widely accepted that the network synchronization is good within each shard [37, 13, 38, 21, 17, 35]. In [37], authors apply a BFT-typed consensus protocol that is suitable for good network condition, yet their protocol still require complex communication. Authors in [2] also propose a consensus protocol for good network condition, but their protocol requires view change to change the leader under malicious cases, reducing efficiency. Some works [33] use Proof-of-Work (PoW) protocol as their intra-shard consensus. However, PoW-based protocol typically has low performance as it cannot guarantee instant finality, and it is easy to fork. In SP-Chain, we exploit the good and synchronous network condition in each shard and propose the two-phase concurrent voting consensus protocol. It reduces the number of communication rounds in consensus (2 rounds) under synchronous network without compromising security, and parallelizes the steps that need to be executed serially in the traditional consensus process, thus significantly improving the consensus efficiency.
In previous sharding systems, the leader is vulnerable to be attacked. The reason is that, in existing protocols [23, 37, 17, 12, 18, 14], a leader keeps producing blocks for a period of time and the rotation of the leader can be known in advance, leaving the chance for attackers to attack the leader. Moreover, when the leader is found to be malicious or attacked, their consensus protocols perform a complex view change process to replace the leader, which is inefficient. In SP-Chain, we design an unbiased distributed randomness generation scheme with low overhead, and propose an efficient and secure leader rotation mechanism based on our consensus protocol and the distributed randomness.
@@
Issues from Cross-Shard Aspect. Cross-shard transaction processing is a significant part of the blockchain sharding system. However, existing cross-shard transaction processing schemes still have drawbacks on protecting security, or they increase large amount of communication overhead when securing transactions. OmniLedger [17] requires an honest client to participate during the cross-shard transaction process. A malicious client can lock the cross-shard transactions and obstruct their executions. Such behavior brings troubles for both system security and efficiency. Some works [37, 13, 12, 18] does not consider the leader to be evil when dealing with cross-shard transactions, making the cross-shard transaction execution less secure. Other works [33, 38] proposes to attach proofs to each cross-shard transaction to protect security. However, this mechanism increases the lots of communication overhead, reducing system efficiency. In SP-Chain, to protect the security of cross-shard transactions with high efficiency, we propose the proof-assisted efficient cross-shard transaction processing mechanism. In this mechanism, pruned Merkle proofs [25] are generated for transactions sent to different shards. With such design, one pruned proof is attached to a batch of transactions to verify them together, reducing the overhead.
2.2 Distributed Randomness
@@
3 System and Threat Model
3.1 System Model
SP-Chain proceeds in epochs. There are multiple slots $t$ in each epoch $e$. We assume there are $n$ nodes in the network for each epoch (noting that $n$ might be changing as epoch changes). Each node is given a public/secret key pair $(PK,SK)$ through a Public-Key Infrastructure (PKI) [38]. All nodes are partitioned into $m$ shards (a.k.a. committees). Thus, there are $k=n/m$ nodes (a.k.a. members) in each shard, including one leader. To prevent Sybil attack, each node is required to generate a Sybil-proof identity when joining the system, leveraging the techniques in [17] and [37].
@@ 3.1 System Model
SP-Chain adopts the account model to represent the state of the blockchain, where each account has its own states. The states of one account are maintained by one certain shard, for computation and storage scalability. Which shard an account’s state should be stored by is determined by its address. The account address is mapped to a shard based on the output of a random oracle (e.g., the remainder of the account address divided by the number of shards). When an account initiates a transaction, that transaction is routed to the corresponding shard based on the address of its sender account. How to design smarter allocation mechanisms for accounts and transactions (e.g., [7, 30, 14, 19]) is orthogonal to this work and will therefore be discussed in our future work.
3.2 Threat Model
We build a similar threat model as previous works do [37, 13, 38]. In our model, there exists a Byzantine adversary who can take control of $<1/3$ fraction of the total nodes. Similar to the previous works, the communication channel is synchronous in one shard. Therefore, each shard can achieve an optimal fault resiliency of 1/2. Corrupted (Byzantine) nodes may collude and behave arbitrarily, such as generating and sending invalid messages (transaction manipulation), sending messages to different nodes with different values (equivocation), or not sending any or all of the messages (silence attack). Other nodes besides the above are called honest nodes, they will always obey the protocol and do not do anything beyond what is specified.
@@ 3.2 Threat Model
4 System Overview
SP-Chain consists of four main components, leader rotation, intra-shard consensus, cross-shard transaction processing, and shard reconfiguration, shown in Fig. 1. Our protocol selects a new leader in each slot to propose one block in each shard based on unbiased randomness. Then, the intra-shard consensus is executed. When the consensus is reached, the (cross-shard) transactions are sent and executed. Each epoch consists of multiple slots followed by a reconfiguration phase, during which the shard members are reshuffled. We now explain each component and the design intuition in more detail.
Leader Rotation. To prevent attacks on the leaders and maintain high performance when leaders are malicious or attacked, we propose the leader rotation scheme. The leaders are changed frequently and randomly to prevent the attackers from knowing the leaders and launching targeted attacks. In each slot, each shard elects a new leader among the nodes. To ensure the security of the leader rotation process, the leader shall be elected based on an unbiased randomness retrieved from the confirmed block. Our distributed randomness generation scheme guarantees high efficiency when electing the new leader, as no extra communications are required. More importantly, as the leaders are rotated automatically in each slot (no matter it is malicious or honest), the view change process is eliminated, achieving high efficiency even under attacks.
Intra-Shard Consensus. To boost the intra-shard consensus performance, we propose the two-phase concurrent voting consensus protocol. We leverage the features of good network synchronization within each shard, and propose an efficient synchronous consensus protocol with less rounds of communication. Moreover, we parallelize the steps that are processed serially in the traditional consensus, and propose the concurrent voting protocol to ensure high block generation and confirmation speed, hence improving system performance.

### v2

@@ Abstract
A promising way to overcome the scalability limitations of the current blockchain is to use sharding, which is to split the transaction processing among multiple, smaller groups of nodes. A well-performing blockchain sharding system requires both high performance and high security in both intra- and cross-shard perspectives. However, existing protocols either have issues in protecting security or trade off great performance for security. In this paper, we propose SP-Chain, a blockchain sharding system with enhanced Security and Performance for both intra- and cross-shard perspectives. For the intra-shard aspect, we design a pipelined two-phase concurrent voting scheme to provide high system throughput and low transaction confirmation latency. Moreover, we propose an efficient unbiased leader rotation scheme to ensure high performance under malicious behavior. For the cross-shard aspect, a proof-assisted efficient cross-shard transaction processing mechanism is proposed to guard cross-shard transactions with low overhead. We implement SP-Chain based on Harmony, and evaluate its performance via large-scale deployment. Extensive evaluations suggest that SP-Chain can process more than 10,000 tx/sec under malicious behaviors with a confirmation latency of 7.6s in a network of 4,000 nodes.
@@
Blockchain, blockchain sharding, cross-shard transaction processing, intra-shard transaction processing
I Introduction
Since the advent of Bitcoin [33], blockchain systems have continued to have a significant impact on society. However, the low system throughput and high transaction confirmation delays of such systems greatly hinder their usability across various infrastructures and applications. Consequently, sharding [7, 30, 48, 2], a promising blockchain scaling solution, has been proposed. Herein, the entire blockchain state is divided into multiple non-overlapping shards, each maintained by a group of nodes.
@@ I Introduction
Intra-shard security and performance: Within each shard, a specific leader is typically selected to propose blocks. Malicious leaders or attacks on leaders can severely affect intra-shard security and performance. Specifically, most existing sharding protocols [30, 21, 50, 12, 22, 23] require each shard’s leader to produce blocks within a certain time frame and to know the next leader in advance. This allows attackers to easily launch targeted attacks on the leaders [25], weakening system security (for example, colluding with the leader for Byzantine behavior or launching DDoS (Distributed Denial-of-Service) attacks against the leader). Moreover, when a leader is found to be malicious or under attack, a complex view change process is required to elect a new leader [5], thereby reducing system performance.
Another issue is that many existing sharding protocols adopt generic Byzantine Fault Tolerance (BFT) consensus protocols for intra-shard consensus, requiring extensive communication among shard members. Specifically, such consensus protocols (e.g., PBFT, Practical Byzantine Fault Tolerance) [5] used in partial-synchronous networks typically require multi-stage communication (e.g., pre-preparation, preparation, commit) to ensure security in poor network connectivity and high latency situations. However, these protocols are inefficient when applied to blockchain sharding. In sharding systems, each shard generally has high network connectivity and a low and fixed message propagation cap [48, 21, 13, 50, 26]. Under such conditions, a more efficient consensus protocol with less communication overhead can be proposed for sharding systems.
@@
Random Leader Rotation. To address the first challenge, the leader of each shard is frequently, randomly, and automatically changed. Specifically, the leaders are rotated on a per-block cycle. To prevent attackers from knowing subsequent leaders in advance, leaders are selected based on distributed randomness only before each block is proposed. To ensure performance and security during the distributed randomness generation process, we propose a chain-based randomness generation mechanism. Herein, nodes within each shard use the signature information of previously confirmed blocks to efficiently generate unbiased distributed randomness. In addition, the automatic leader rotation mechanism eliminates the originally complex view change process, further improving the efficiency of leader election.
@@
More importantly, to further enhance consensus efficiency, we make both the computation and communication steps run in parallel, whereas in traditional consensus protocols these steps are processed serially. In traditional protocols, the leader first packages transactions into a block. Then, the leader broadcasts the block. Nodes receiving the block will vote and reach a consensus. After reaching a consensus, nodes commit the block into the blockchain. Our concurrent voting protocol parallelizes these computation and communication steps. Specifically, when a shard’s leader produces a new block, the members of that shard vote on and reach a consensus on the previous block. Also, while the leader broadcasts the new block, each node commits the previous block into the blockchain. This proposed scheme provides fast block generation and confirmation rates, bringing high throughput and low latency to the system.
@@
This paper mainly makes the following contributions:
  • We propose SP-Chain, a novel blockchain sharding system that significantly improves both intra-shard and cross-shard security and performance.
  • We design a pipelined two-phase concurrent voting protocol for intra-shard consensus, which greatly increases throughput and lowers latency compared to traditional sequential BFT consensus.
  • We introduce an unbiased leader rotation mechanism based on signature-derived randomness that ensures no predictable leader pattern, enhancing security against leader targeting attacks while maintaining high efficiency.
  • We develop an efficient proof-assisted cross-shard transaction processing scheme that securely handles transactions across shards with minor overhead.
  • We implement a prototype of SP-Chain based on Harmony [^1] A well-known public blockchain sharding project that was once ranked in the top 50 in the cryptocurrency space in terms of market capitalization [42] and conduct extensive large-scale experiments (up to 4,000 nodes). Our evaluation demonstrates that SP-Chain achieves over 10,000 transactions per second (TPS) with low confirmation latency, outperforming state-of-the-art sharding systems.
II Background and Related Work
II-A Blockchain Sharding
@@
A well-performing blockchain sharding system needs to ensure good security as well as high performance. Unlike traditional blockchain systems, in blockchain sharding systems, there are a large number of cross-shard transactions that are transmitted among shards [44, 21, 48, 35, 14, 24, 36]. Nodes need to not only process transactions within a shard, but also handle transactions across shards. Therefore, it is necessary to consider both intra- and cross-shard aspects when analyzing the security and performance of blockchain sharding. However, previous works have problems with both intra- and cross-shard security and performance.
@@
Issues from Cross-Shard Aspect. Cross-shard transaction processing is a significant part of the blockchain sharding system. However, existing cross-shard transaction processing schemes still have drawbacks on protecting security, or they increase large amount of communication overhead when securing transactions. OmniLedger [21] requires an honest client to participate during the cross-shard transaction process. A malicious client can lock the cross-shard transactions and obstruct their executions. Such behavior brings troubles for both system security and efficiency. Some works [48, 13, 12, 22] does not consider the leader to be evil when dealing with cross-shard transactions, making the cross-shard transaction execution less secure. Other works [44, 50] proposes to attach proofs to each cross-shard transaction to protect security. However, this mechanism increases the lots of communication overhead, reducing system efficiency. BrokerChain [15] introduces special "broker" accounts to reduce cross-shard transfers in account-based sharding, helping balance the workload across shards. However, its intermediary accounts add protocol complexity. Authors in [29] propose a flexible sharding protocol that integrates a cross-shard Byzantine fault tolerance (CSBFT) consensus, allowing multiple shards to coordinate on cross-shard transactions in a single consensus round and thereby cutting down confirmation delay. The CSBFT, however, incurs extra cross-shard coordination overhead for each transaction. LightCross [38] focuses on minimizing cross-shard communication rounds by using trusted hardware (TEE). Nonetheless, LightCross’s reliance on TEE devices introduces additional trust and hardware assumptions. ShardCon [49] embeds cross-shard transaction handling into a unified Byzantine consensus spanning the involved shards. However, coordinating a BFT consensus among multiple shards for each cross-shard transaction incurs high communication and computation costs, especially as the number of shards grows.
In SP-Chain, to protect the security of cross-shard transactions with high efficiency, we propose the proof-assisted efficient cross-shard transaction processing mechanism. Instead of heavy cross-shard proofs, special hardware reliance, complex two-phase commit exchanges, or full multi-shard consensus for each transaction, SP-Chain attaches compact cryptographic proofs (e.g., pruned Merkle proofs [28]) to cross-shard transactions, allowing involved shards to independently verify transaction outcomes. With such design, one pruned proof is attached to a batch of transactions to verify them together, reducing the overhead. This design secures cross-shard transactions (even against adversarial senders or leaders) with high efficiency.
Issues from Intra-Shard Aspect. Existing sharding works reach consensus with limited efficiency. A major reason is that most of them [30, 21, 13, 8, 50, 12, 14, 22, 29, 23] use BFT-typed consensus protocols suitable for scenarios with poor network conditions (i.e., partially synchronous or asynchronous). Those protocols usually require multiple rounds of communication ($\geq$3) to reach consensus, slowing down the efficiency of reaching consensus. However, it is widely assumed that the network has reasonably good synchronization within each shard (i.e., bounded delays), as is common in such systems [48, 13, 50, 26, 21]. In [48], authors apply a BFT-typed consensus protocol that is suitable for good network condition, yet their protocol still require complex communication. Authors in [1] also propose a consensus protocol for good network condition, but their protocol requires view change to change the leader under malicious cases, reducing efficiency. Some works [44, 24] use Proof-of-Work (PoW) protocol as their intra-shard consensus. However, PoW-based protocol typically has low performance as it cannot guarantee instant finality, and it is easy to fork. Cherubim [27] introduces a quadruple-pipelined two-phase commit (4P-2PC) that parallelizes intra-shard consensus and cross-shard transaction processing to boost throughput. However, the scheme’s complexity still incurs significant coordination overhead. In SP-Chain, we exploit the good and synchronous network condition in each shard and propose the two-phase concurrent voting consensus protocol. It reduces the number of communication rounds in consensus (2 rounds) under synchronous network without compromising security, and parallelizes both the computation and communication steps that would otherwise be handled serially in the traditional consensus process, thus significantly improving the consensus efficiency.
In previous sharding systems, the leader is vulnerable to be attacked. The reason is that, in existing protocols [30, 48, 21, 12, 22, 14], a leader keeps producing blocks for a period of time and the rotation of the leader can be known in advance, leaving the chance for attackers to attack the leader. Moreover, when the leader is found to be malicious or attacked, their consensus protocols perform a complex view change process to replace the leader, which is inefficient. Despite the emergence of various new types of blockchain sharding systems in the last few years, most of them, however, also suffer from the aforementioned vulnerability of the leader to attacks as well as malicious leaders affecting performance [24, 23, 29, 27, 38, 49, 15, 46]. For example, Cherubim [27] treats the underlying Byzantine consensus as a black box and does not specifically address leader fairness – a malicious or monopolizing leader could still hinder performance or bias the process. LightCross [38] does not modify the intra-shard consensus (it builds on an existing PBFT-style engine FISCO-BCOS), so issues of leader fairness or suboptimal consensus throughput per shard remain unaddressed. ShardCon’s [49] design prioritizes cross-shard consistency and formal security proofs, but it does not fundamentally enhance intra-shard consensus performance or address the potential performance bottleneck of a faulty leader before a view change triggers. In SP-Chain, we design an unbiased distributed randomness generation scheme with low overhead, and propose an efficient and secure leader rotation mechanism that regularly and randomly rotates the shard leader, based on our consensus protocol and the distributed randomness. Consequently, SP-Chain avoids the single-leader performance bottleneck and mitigates malicious leader behavior, a feature absent in most of the recent state-of-the-art blockchain sharding systems (e.g., [29, 27, 38, 49, 15, 46, 18, 17]).
II-B Distributed Randomness
@@
II-C Alternative Cross-Shard Transaction Processing Schemes
Alternatively, cross-shard transaction execution can be offloaded to layer-2 mechanisms such as state channels [34] and Plasma chains [37]. In a state-channel-based approach, participants lock assets on-chain and conduct numerous transfers off-chain, only settling the final state on-chain. This avoids most on-chain cross-shard communication but requires continuous client involvement – users must remain online (or employ trusted "watchers") to respond to disputes or malicious counter-parties, and the method typically only supports a limited set of participants per channel. Plasma sidechains take a different route by anchoring sharded (or child-chain) transactions to a main chain via periodic commitments. Plasma designs can facilitate many off-chain transactions, but they introduce significant trust assumptions and latency: users depend on an honest Plasma operator and must engage in complex exit protocols to withdraw funds safely. For example, funds withdrawals from a Plasma chain are often delayed by a predetermined challenge period (on the order of one to two weeks) to allow fraud proofs, and users need to actively monitor the system to ensure the security of their assets. These limitations – high client burden, delayed finality, and reliance on external trusted parties or timely responses – make purely off-chain approaches less desirable for cross-shard verification in an open system.
In contrast, SP-Chain addresses the above limitations within its on-chain sharding design. The cross-shard transaction mechanism of SP-Chain is integrated into the protocol so that no special action is required from clients beyond initiating the transaction. Unlike OmniLedger’s client-dependent cross-shard commits or state channels that assume 100% availability of participants, SP-Chain achieves atomic cross-shard confirmation without user intervention. This is enabled by a proof-assisted verification: each cross-shard transaction carries a cryptographic proof (using pruned Merkle proofs) that can be batched and verified by the destination shard’s validators, eliminating any need to trust a separate intermediary (as in Plasma) or to exchange multiple round-trip messages off-chain. By enforcing cross-shard validity through the consensus of involved shards, SP-Chain maintains security under standard BFT assumptions and avoids introducing new trust dependencies. In summary, SP-Chain’s approach preserves the trustless security of on-chain sharding while minimizing overhead, effectively overcoming the client participation requirements and security vulnerabilities of state-channel and Plasma-based solutions.
III System and Threat Model
III-A System Model
[…]
@@ III-A System Model
SP-Chain adopts the account model to represent the state of the blockchain, where each account has its own states. The states of one account are maintained by one certain shard, for computation and storage scalability. Which shard an account’s state should be stored by is determined by its address. The account address is mapped to a shard based on the output of a random oracle (e.g., the remainder of the account address divided by the number of shards). When an account initiates a transaction, that transaction is routed to the corresponding shard based on the address of its sender account. How to design smarter allocation mechanisms for accounts and transactions (e.g., [6, 39, 14, 24, 38]) is orthogonal to this work and will therefore be discussed in our future work.
III-B Threat Model
We build a similar threat model as previous works do [48, 13, 50]. In our model, there exists a Byzantine adversary who can take control of <1/3<1/3< 1 / 3 fraction of the total nodes. Similar to the previous works, the communication channel is synchronous in one shard. Therefore, each shard can achieve an optimal fault resiliency of 1/2. Corrupted (Byzantine) nodes may collude and behave arbitrarily, such as generating and sending invalid messages (transaction manipulation), sending messages to different nodes with different values (equivocation), or not sending any or all of the messages (silence attack). Other nodes besides the above are called honest nodes, they will always obey the protocol and do not do anything beyond what is specified.
@@ III-B Threat Model
To mitigate Sybil attacks, similar to existing works [21, 48, 13, 24, 42], each participating node must obtain a Sybil-resistant identity through an offline proof-of-work puzzle, or proof-of-stake, or PKI-based identity issuance process. This requirement makes the generation of new identities expensive and verifiable, preventing a malicious adversary from cheaply acquiring arbitrarily many node identities. Thus, the adversary is effectively limited to controlling at most the assumed fraction (<1/3) of all nodes. Moreover, at each epoch’s shard reconfiguration, node identities are uniformly re-distributed across shards and validated (see Section V-D), ensuring that no shard can be overwhelmed by colluding Sybil nodes. These measures, built into the protocol, collectively enforce Sybil-resistance under our adversary model.
IV System Overview
SP-Chain consists of four main components, leader rotation, intra-shard consensus, cross-shard transaction processing, and shard reconfiguration, shown in Figure 1. Our protocol selects a new leader in each slot to propose one block in each shard based on unbiased randomness. Then, the intra-shard consensus is executed. When the consensus is reached, the (cross-shard) transactions are sent and executed. Each epoch consists of multiple slots followed by a reconfiguration phase, during which the shard members are reshuffled. We now explain each component and the design intuition in more detail.
Leader Rotation. To prevent attacks on the leaders and maintain high performance when leaders are malicious or attacked, we propose a leader rotation scheme. The leader rotation in SP-Chain is a mechanism by which a new block proposer is chosen every block-producing cycle (i.e., one slot). Specifically, the leaders are changed frequently and randomly to prevent the attackers from knowing the leaders and launching targeted attacks. In each slot, each shard elects a new leader among the nodes. To ensure the security of the leader rotation process, the leader shall be elected based on an unbiased randomness retrieved from the confirmed block. Our distributed randomness generation scheme guarantees high efficiency when electing the new leader, as no extra communications are required. More importantly, as the leaders are rotated automatically in each slot (no matter it is malicious or honest), the view change process is eliminated, achieving high efficiency even under attacks.
Intra-Shard Consensus. To boost the intra-shard consensus performance, we propose the two-phase concurrent voting consensus protocol. We leverage the features of good network synchronization within each shard, and propose an efficient synchronous consensus protocol with less rounds of communication. More importantly, we make both computation and communication steps run in parallel (which are handled sequentially in previous protocols), and propose the concurrent voting protocol to ensure high block generation and confirmation speed, hence improving system performance.